# Doped resonating valence bond states: How robust are the spin ice phases in 3D Rydberg arrays

### v1

@@ Abstract
Rydberg blockade effect provides a convenient platform for simulating locally constrained many-body systems, such as quantum dimer models and quantum loop models, especially their novel phases like topological orders and gapless quantum spin ice (QSI) phases. To discuss the possible phase diagram containing different QSIs in 3D Rydberg arrays, here, we have constructed an extended Rokhsar-Kivelson (RK) Hamiltonian with equal-weight-superposition ground state in different fillings at the RK point. Therefore, both the perfect QSIs with fixed local dimer filling and their monomer-doped states can be simulated directly by Monte Carlo sampling. Using single mode approximation, the excitations of dimers and monomers have also been explored in different fillings. We find that, in the thermodynamical limit, even doping a small amount of monomers can disrupt the topological structure and lead to the existence of off-diagonal long-range order. However, in a finite size (as in cold-atom experiment), the property of QSI will be kept in a certain region like a crossover after doping. The phase diagram containing different QSIs and off-diagonal order phases is proposed.
@@ I Introduction
In the past years, programmable Rydberg atom arrays have played the roles of powerful quantum simulators for the investigation of highly correlated and entangled quantum matter [1, 2, 3, 4, 5]. These systems have opened up new avenues to study novel quantum many-body phases [6, 7, 8, 9, 10], complex dynamics [11, 12, 13], gauge theories [14, 15, 16, 17], and even combinatorial optimization problems [18, 19, 20].
@@
An especially promising direction that has recently attracted much attention is the simulation of quantum phases of matter in these tunable atomic setups. Such phases and the transitions between them have been intensely studied for Rydberg atoms arrayed in one spatial dimension [21, 22, 23, 24] as well as in various two-dimensional geometries, including square [25, 26, 27, 28, 29], triangular [30, 31], honeycomb [32, 33], kagome [34, 35], and ruby lattices [36, 37, 38]. In particular, Ref. [34] identified an intriguing highly correlated regime in the phase diagram of the kagome-lattice Rydberg atom array, as characterized by a lack of symmetry-breaking solid order and a large entanglement entropy. The correlations in this region were found to be “liquid-like”, in that the Rydberg excitations are highly correlated and constrained by strong Rydberg-Rydberg repulsions, as opposed to a weakly interacting gas where atomic excitations at different sites fluctuate almost individually. Such a correlated liquid phase can be understood if we map this system onto a quantum dimer model (QDM) on the triangular lattice [39, 40, 41, 42, 43, 44]. Under this mapping, the Rydberg blockade effect can dynamically exert certain local constraints, such that each dual triangular lattice site is attached to some fixed number of dimers. On the non-bipartite triangular lattice, this local constraint further enables mapping to the $\mathbb{Z}_{2}$ lattice gauge theory, which raises the possibility of liquid phases with $\mathbb{Z}_{2}$ topological order [34, 42, 35]. Moreover, the dimer density can be tuned continuously by the laser detuning, which acts as an effective chemical potential of dimers. For the example of the kagome lattice, dimer filling from one per site to two per site has been found in numerics [34, 35, 10].
@@
On the other hand, Rydberg experiments have also been extended to investigations of QDM physics on bipartite lattices. In contrast to the cases non-bipartite lattices that host $\mathbb{Z}_{2}$ gauge structures, QDMs on bipartite lattices generally host a continuous $U(1)$ gauge structure. In 3D bipartite lattices, QDMs [45, 39] have long been known to host the so-called Coulomb phases that exhibit deconfined charges and gapless “photon” excitations, just analogous to the Maxwell’s theory of electromagnetism [46, 47, 48, 49, 50]. Such a $U(1)$ Coulomb phase has been extensively studied in the context pyrochlore lattice frustrated quantum magnets, and were named the $U(1)$ quantum spin ice (QSI) in that context. In two spatial dimension, however, the “photons” will be gapped out due to the Polyakov’s confinement mechanism [39, 51, 52, 53, 31], hence the deconfined $U(1)$ Coulomb phase cannot exist as a stable phase in 2D. As such, experimental realization of the Coulomb QSI phase requires Rydberg simulators in three spatial dimension [54, 55, 56].
@@
In this work, we explore a crucial yet previously unexplored extension of the QDM on 3D bipartite lattices [46, 47, 57, 39]. Unlike the original QDMs, our model allows for a variable dimer density in which monomer doping is incorporated. For original QDMs, the local constraint that each site is attached to a given number of dimers is crucial for the emergent gauge structure. However, in Rydberg experiments, monomer defects violating such local constraint can be presented for various reasons [9]. In the previous study, we have examined the case of triangular lattice QDM [35], which relates to the case where atoms are positioned on the sites of the kagome lattice [41, 58, 34]. This configuration is experimentally accessible, and initial investigations of such a system have already been conducted by the team in Ref. [9]. In this work, we turn our attention to 3D bipartite lattices with emergent $U(1)$ gauge structure, and discover the physics with monomer doping present.
@@
Besides the inevitable presence of monomers in experiments, the robustness of $U(1)$ Coulomb liquid phase in presence of monomers itself is an issue that worth further numerical explorations. Here we show that with monomer doping, the 3D QDM exhibits novel features relevant to ongoing Rydberg-atom experiments. More interestingly, the phase diagram of this extended QDM also reveals several distinct $U(1)$ Coulomb liquids with different dimer fillings on each site. Note that in early discussions of such $U(1)$ Coulomb liquids in dimer models, the different phases are characterized by their background charges, or, the number of dimers attached to each site [56, 55, 59, 60]. In the present model, the number of dimers on each site fluctuates among different values; nevertheless, the distinction of the different QSIs still survives based on the different local charges of the background [61]. However, in the case of a soft constraint as is relevant here, we will investigate this fate numerically with Monte Carlo (MC) simulations.
@@
In light of the situation, here, we overcome this challenge and present large-scale Monte Carlo simulation results on the RK-like Hamiltonian with tunable doping in diamond and cubic lattice QDMs. Surprisingly, while $U$(1) Coulomb liquids can be stable in 3D, our numerical results for large system sizes – in both static and dynamical properties – reveal the fragility of $U$(1) Coulomb phases upon doping. Nevertheless, the good news is that, at smaller length scales, the properties of QSIs are preserved under weak doping. This suggests that achieving $U(1)$ Coulomb liquids in cold-atom experiments, which exhibit limited system sizes, remains feasible even in presence of moderate doping.
@@
The $U(1)$ Coulomb liquid phase on a 3D pyrochlore lattice has been studied in a long time both in condensed matter and cold-atom systems [62, 56, 59, 57, 63, 55]. This phase is also referred to as “$U(1)$ QSI” in the context of pyrochlore quantum Ising magnets, and can be well understood in terms of the QDM on the dual diamond lattice. The geometry of the pyrochlore lattice is shown in Fig. 1 and 2, where each pyrochlore site is shared by one up-pointing tetrahedron and one down-pointing tetrahedron. The center of these tetrahedra form a bipartite diamond lattice. For convenience in later discussions, we label the centers of the up-pointing and down-pointing tetrahedra as A and B sublattices of the diamond lattice, respectively. Before discussing the low-energy effective model of the Rydberg arrays, let us review the realistic Hamiltonian in Rydberg experiments:
[…]
@@
[…]
@@
with $h=\delta-\frac{1}{2N}\sum_{i\neq j}V_{ij}$.
Notice that the van der Waals interaction $V_{ij}$ term leads to the blockade effect, meaning that Rydberg excitations within the blockade radius $R_{b}$ experience a large pairwise repulsive interactions hence should be suppressed. The consequence of such Rydberg blockade effect is clearly seen if we only keep the van der Waals interaction at the nearest neighbor $V_{1}$, where further neighbor interactions are ignored due to the rapid $\sim r^{-6}$ decay of the van der Waals potential. $H_{0}$ is then expressed as:
@@
For the limit $V_{1}\gg\Omega$, $\hat{H}_{\Omega}$ can be regarded as a perturbation. The low-energy effective Hamiltonian under Schrieffer-Wolff formulation of perturbation theory up to the sixth order can be obtained [55]. Boson creation or annihilation at a single site violates the ice rule, so the first-order contribution clearly vanishes. A similar situation arises at all odd orders, so we only consider corrections at even orders. At second order, bosons must be created and annihilated at a single site to return to the ice manifold. This can occur at any point, contributing a constant term to the energy. At fourth order, bosons hop along a link to a neighboring site and then return along the same link, contributing a constant energy. Such processes are present for all links. At sixth order, one process involves a single boson hopping around the triangular face of each tetrahedron. Another involves alternating ground and excited atoms on a hexagonal cell as shown in Fig. 2 (b), which can be mapped to flippable configurations through a ring-exchange process. The former contributes a constant energy. Ignoring all these constant terms, the effective Hamiltonian is a six-sites ring exchange term
$H_{\textrm{eff}}^{(6)}=-J_{ring}\sum\left(S_{1}^{+}S_{2}^{-}S_{3}^{+}S_{4}^{-}S_{5}^{+}S_{6}^{-}+h.c.\right),$ (4)
@@
where the sum is over all the hexagonal cells in the pyrochlore lattices. $J_{ring}$ can be obtained by summing over virtual processes and $S_{i}^{\pm}$ represent that boson (spin) creation or annihilation at a single site $i$.
@@
For mapping to a QDM on the dual lattice, we connect the centers of all nearest-neighbour tetrahedra as shown in Fig. 1 (b) and 2 (b), with each atom exactly at the center of each link. Atoms in the Rydberg state (spin up) can be mapped to dimers in the dual (diamond) lattice, while atoms in the ground state (spin down) can be mapped to empty links. The high-order ring exchange dynamics becomes the effective kinetic term. Thus the QDM on diamond lattice takes the role of low-energy effective description of this pyrochlore Rydberg array [55]. The ground state of quantum dimer models can be exactly solved by adding a potential energy term $V$ and adjusting it to a particular value $V=J$, referred to as the Rokhsar-Kivelson (RK) point [39].
@@
where the sum is performed over all hexagonal plaquettes. The first term is ring exchange term and the second term is potential energy term.
@@
When $t/V=1$, i.e., at the famous RK point [39], the ground state is a resonating valence bond (RVB) state described by the coherent equal-weight superposition of all the dimer covering configurations, i.e., $|G\rangle=\sum_{C}|C\rangle$. Moreover, the RK point exhibit topological degeneracy, where the equal-weight superpositions within different topological sectors are ground states with identical energies [64].
@@
Figure 1: The mapping between spin, dimer and flux. (a) Rydberg array can be mapped to spin model. The atom at the Rydberg excited state is mapped to the spin-up (up arrows), while the atom at the atomic ground state is mapped to the spin-down (down arrows).(b) Spins can be mapped onto dimer (orange line), A dimer, corresponds to a spin pointing upward, $S_{z}=1/2$, while the absence of a dimer (blue line) corresponds to a spin pointing downward, $S_{z}=-1/2$. (c) They can be interpreted as fluxes, we identify a spin pointing up, with a flux vector $\mathbf{B}$ pointing inward from a tetrahedron, and vice versa. The ice rule corresponds to contact two upward spins, two dimers and the flux field vanishes for each tetrahedron.
@@
In the phase diagram of QDMs on 3D bipartite lattices, the RK point is located as a part of the $U(1)$ Coulomb phase. In the spin language, the local constraint of dimers corresponds to the ice rule that the total spin in each tetrahedron of the pyrochlore lattice is same. In the dimer language, it means that the number of dimers attached to each site of the dual diamond lattice is fixed. Further, this local constraint is an essential ingredient that allows a mapping to the $U(1)$ gauge theory: Spin-up (dimer) and spin-down (empty link) can be mapped to the ±$\pm\frac{1}{2}$ magnetic flux pointing from A to B diamond sublattices, as displayed in Fig. 1. This leads to the condition $\left(\mathrm{div}B\right)_{\mathsf{r}}=0$ for the case of two-dimer-per-site configurations, and $\left(\mathrm{div}B\right)_{\mathsf{r}}=\eta_{\mathsf{r}}$ for the one-dimer-per-site configurations. Here $\mathsf{r}$ denotes the site of the dual diamond lattice, and $\eta_{\mathsf{r}}=\pm 1$ for A and B sublattices, respectively. Such divergence condition is analogous to the Gauss’s law of electromagnetism, and the background charges can be absorbed by redefining the strength of magnetic field.
@@
The pyrochlore lattice is a composite lattice, consisting of four sets of face-centered cubic lattices. As shown in Fig. 2 (c), we label the points on the tetrahedron belonging to different sublattices as 1, 2, 3, and 4. In Fig. 2 (d), we display one of the sublattices and define its direct lattice vectors as $\overrightarrow{a_{1}},\overrightarrow{a_{2}},\overrightarrow{a_{3}}$, with the corresponding reciprocal lattice vectors as $\overrightarrow{b_{1}},\overrightarrow{b_{2}},\overrightarrow{b_{3}}$ (see Fig. 2 (e)).

### v2

# Doped resonating-valence-bond states: Robustness of the spin-ice phases in three-dimensional Rydberg arrays

@@ Abstract
Rydberg blockade effect provides a convenient platform for simulating locally constrained many-body systems, such as quantum dimer models and quantum loop models, especially their novel phases like topological orders and gapless quantum spin ice (QSI) phases. To discuss the possible phase diagram containing different QSIs in three-dimensional (3D) Rydberg arrays, we have constructed an extended Rokhsar-Kivelson (RK) Hamiltonian with equal weight superposition ground state in different fillings at the RK point. Therefore, the perfect QSIs with fixed local dimer filling and their monomer-doped states can be simulated directly by Monte Carlo sampling. Using single-mode approximation, the excitations of dimers and monomers have also been explored in different fillings. We find that, in the thermodynamical limit, even doping a small amount of monomers can disrupt the topological structure and lead to the existence of off-diagonal long-range order. However, in a finite size (as in cold-atom experiment), the property of QSI will be kept in a certain region like a crossover after doping. The phase diagram containing different QSIs and off-diagonal order phases is proposed.
@@ I Introduction
In recent years, programmable Rydberg atom arrays have played the roles of powerful quantum simulators for the investigation of highly correlated and entangled quantum matter [1, 2, 3, 4, 5, 6]. These systems have opened up new avenues to study novel quantum many-body phases [7, 8, 9, 10, 11], complex dynamics [12, 13, 14], gauge theories [15, 16, 17, 18], and even combinatorial optimization problems [19, 20, 21].
@@
An especially promising direction that has recently attracted much attention is the simulation of quantum phases of matter in these tunable atomic setups. Such phases and the transitions between them have been intensely studied for Rydberg atoms arrayed in one spatial dimension [22, 23, 24, 25] as well as in various two-dimensional (2D) geometries, including square [26, 27, 28, 29, 30], triangular [31, 32], honeycomb [33, 34], kagome [35, 36], and ruby lattices [37, 38, 39]. In particular, Ref. [35] identified an intriguing highly correlated regime in the phase diagram of the kagome-lattice Rydberg atom array, characterized by a lack of symmetry-breaking solid order and a large entanglement entropy. The correlations in this region were found to be “liquid-like”, in that the Rydberg excitations are highly correlated and constrained by strong Rydberg-Rydberg repulsions, as opposed to a weakly interacting gas where atomic excitations at different sites fluctuate almost individually. Such a correlated liquid phase can be understood if we map this system onto a quantum dimer model (QDM) on the triangular lattice [40, 41, 42, 43, 44, 45]. Under this mapping, the Rydberg blockade effect can dynamically exert certain local constraints, such that each dual triangular lattice site is attached to some fixed number of dimers. On the non-bipartite triangular lattice, this local constraint further enables mapping to the $\mathbb{Z}_{2}$ lattice gauge theory, which raises the possibility of liquid phases with $\mathbb{Z}_{2}$ topological order [35, 43, 36]. Moreover, the dimer density can be tuned continuously by laser detuning, which acts as an effective chemical potential of dimers. For the example of the kagome lattice, dimer filling from one per site to two per site has been found in numerics [35, 36, 11].
@@
On the other hand, Rydberg experiments have also been extended to investigations of QDM physics on bipartite lattices. In contrast to the cases of non-bipartite lattices that host $\mathbb{Z}_{2}$ gauge structures, QDMs on bipartite lattices generally host a continuous U(1) gauge structure. In three-dimensional (3D) bipartite lattices, QDMs [46, 40] have long been known to host the so-called Coulomb phases that exhibit deconfined charges and gapless “photon” excitations, just analogous to Maxwell’s theory of electromagnetism [47, 48, 49, 50, 51]. Such a U(1) Coulomb phase has been extensively studied in the context of pyrochlore lattice frustrated quantum magnets and was named the U(1) quantum spin ice (QSI) in that context. In two spatial dimensions, however, the “photons” will be gapped out due to the Polyakov’s confinement mechanism [40, 52, 53, 54, 32], hence the deconfined U(1) Coulomb phase cannot exist as a stable phase in 2D. As such, the experimental realization of the Coulomb QSI phase requires Rydberg simulators in three spatial dimensions [55, 56].
In recent years, with the development of quantum simulation technologies, platforms such as Rydberg atom arrays have provided new avenues for the nonequilibrium preparation of topological quantum states. Notably, the topological features observed on these platforms may not originate from the equilibrium ground state of the system. For instance, Ref. [10] reported spin-liquid–like behavior that was in fact realized through nonequilibrium dynamical processes. Subsequent theoretical studies have revealed the underlying mechanism: the response rates of e-anyons and m-anyons to external parameter variations differ significantly—while e-anyons can adiabatically follow the parameter changes and become annihilated, m-anyons are preserved due to dynamical freezing. As a result, a quantum spin liquid with few defects emerges at the macroscopic scale Ref. [57]. This mechanism has not only been confirmed by extensive numerical simulations but also exhibits strong universality, thereby opening new directions for the preparation and stabilization of topological states in finite systems.
The recent Ref.[58] pointed out that the low-energy effective model of the pyrochlore Rydberg array is the diamond quantum dimer model. They give an experimental scheme to realize the quantum spin ice phases based on this mapping between the dimer model and the Rydberg array. The effective model we adopt is highly related to this experimental proposal.
In this work, we explore a crucial yet previously unexplored extension of the QDM on 3D bipartite lattices [47, 48, 59, 40]. The main aim of our work is to explore how robust the spin ice in this system, thus the quantum spin ice of Rydberg array can be effectively considered as the resonating valence bond (RVB) in 3D quantum dimer models. Under realistic experimental conditions, the local constraint (Rydberg blockade) cannot be perfectly enforced, and as a result, some monomers usually remain. Therefore, the key question is, do the spin ice still exist after doping in RVB, and are there some remaining characteristics of spin ice in a finite size system even if the spin ice disappear? Our numerical results actually have answered these questions generally, not limited to the detailed Rydberg arrays.
@@
Unlike the original QDMs, our model allows for a variable dimer density in which monomer doping is incorporated. For original QDMs, the local constraint that each site is attached to a given number of dimers is crucial for the emergent gauge structure. However, in Rydberg experiments, monomer defects violating such local constraint can be presented for various reasons [10]. In the previous study, we examined the case of triangular lattice QDM [36], which relates to the case where atoms are positioned on the sites of the kagome lattice [42, 60, 35]. The QDM is experimentally accessible and has already been conducted initial investigations in Ref. [10], where a dimer model on the kagome lattice. In this work, we turn our attention to 3D bipartite lattices with emergent U(1) gauge structure, and discover the physics with monomer doping present.
@@
Besides the inevitable presence of monomers in experiments, the robustness of U(1) Coulomb liquid phase in presence of monomers itself is an issue that is worth further numerical explorations. Here we show that with monomer doping, the 3D QDM exhibits novel features relevant to ongoing Rydberg-atom experiments. More interestingly, the phase diagram of this extended QDM also reveals several distinct U(1) Coulomb liquids with different dimer fillings on each site. Note that in early discussions of such U(1) Coulomb liquids in dimer models, the different phases are characterized by their background charges, or, the number of dimers attached to each site [56, 58, 61, 62]. In the present model, the number of dimers on each site fluctuates among different values; nevertheless, the distinction of the different QSIs still survives based on the different local charges of the background [63]. However, in the case of a soft constraint as is relevant here, we will investigate this fate numerically with Monte Carlo simulations.
@@
In light of the situation, here, we overcome this challenge and present large-scale Monte Carlo simulation results on the Rokhsar-Kivelson (RK) like Hamiltonian with tunable doping in diamond and cubic lattice QDMs. Surprisingly, while U(1) Coulomb liquids can be stable in 3D, our numerical results for large system sizes – in both static and dynamical properties – reveal the fragility of U(1) Coulomb phases upon doping. Nevertheless, the good news is that, at smaller length scales, the properties of QSIs are preserved under weak doping. Considering that current Rydberg atom array platforms are inherently finite-sized systems, the features of the U(1) Coulomb liquids remain visible at low monomer doping densities. Therefore, we believe that Rydberg atom experiments offer a promising platform for detecting the property of U(1) Coulomb liquids.
@@
The U(1) Coulomb liquid phase on a 3D pyrochlore lattice has been studied in a long time both in condensed matter and cold-atom systems [64, 56, 61, 59, 65, 58]. This phase is also referred to as “U(1) QSI” in the context of pyrochlore quantum Ising magnets, and can be well understood in terms of the QDM on the dual diamond lattice. The geometry of the pyrochlore lattice is shown in Fig. 1 and Fig. 2, where each pyrochlore site is shared by one up-pointing tetrahedron and one down-pointing tetrahedron. The centers of these tetrahedra form a bipartite diamond lattice. For convenience in later discussions, we label the centers of the up-pointing and down-pointing tetrahedra as a and b sublattices of the diamond lattice, respectively. Before discussing the low-energy effective model of the Rydberg arrays, let us review the realistic Hamiltonian in Rydberg experiments:
$\hat{H}$= $\sum_{i=1}^{N}\left[\frac{\Omega}{2}\left(\left|g\right>_{i}\left<r\right|_{i}+\left|r\right>_{i}\left<g\right|_{i}\right)-\delta\left|r\right>_{i}\left<r\right|_{i}\right]$
$+\sum_{i,j=1}^{N}\frac{V_{ij}}{2}\left(\left|r\right>_{i}\left<r\right|_{i}\otimes\left|r\right>_{j}\left<r\right|_{j}\right),$ (1)
@@
The above Rydberg model can be mapped to the quantum Ising model with both transverse and longitudinal fields. We can map the two levels of atoms to effective spin-$1/2$ objects: $\frac{1}{2}\left(\left|g\right>_{i}\left<r\right|_{i}+\left|r\right>_{i}\left<g\right|_{i}\right)=\hat{S}_{i}^{x}$ and $\left|r\right>_{i}\left<r\right|_{i}=\hat{n}_{i}=\hat{S}_{i}^{z}+\frac{1}{2}$. The Hamiltonian takes the same form $H=\hat{H}_{0}+\hat{H}_{\Omega}$ in studies of finite-size systems and thermodynamic limit [8, 26, 66], where
@@
with effective local longitudinal field $h=\delta-\frac{1}{2N}\sum_{i\neq j}V_{ij}$. The local field can induce single-spin flips, thereby altering the density of atoms excited to the Rydberg state. In studies of finite-size systems, it is essential to account for the impact of such local effects. To this end, we have examined cases with different filling densities in our subsequent analysis, specifically: only one atom within each tetrahedron is excited to the Rydberg state, and two atoms within each tetrahedron are excited to the Rydberg state.
Notice that the van der Waals interaction $V_{ij}$ term leads to the blockade effect, meaning that Rydberg excitations within the blockade radius $R_{b}$ experience large pairwise repulsive interactions and hence should be suppressed. The blockade radius is given by $R_{b}=(C_{6}/\hbar\Omega)^{1/6}$, where and $\hbar$ is the reduced Planck constant. Specifically, when the blockade radius is equal to the distance between nearest-neighbor atoms in the pyrochlore lattice, the lattice strictly satisfies the local constraint (at most one atom per tetrahedron can be excited to the Rydberg state). The consequence of such Rydberg blockade effect is clearly seen if we only keep the van der Waals interaction at the nearest neighbor $V_{1}$, where further neighbor interactions are ignored due to the rapid $\sim r^{-6}$ decay of the van der Waals potential. $H_{0}$ is then expressed as:
@@
For the limit $V_{1}\gg\Omega$, $\hat{H}_{\Omega}$ can be regarded as a perturbation. The low-energy effective Hamiltonian under Schrieffer-Wolff formulation of perturbation theory up to the sixth order can be obtained [58]. We will describe the perturbative process using the language of hard-core boson, which will hereafter be referred to simply as ’boson’. To simplify the discussion, we derive here the effective Hamiltonian for the case of n=2 (i.e., two dimers per site). Since the creation or annihilation of a boson on a single lattice site would violate the ice rule, it would lead to the appearance of either three dimers per site or one dimer per site. So the first-order contribution clearly vanishes. A similar situation arises at all odd orders, so we only consider corrections at even orders. At the second order, bosons must be created and annihilated at a single site to return to the ice manifold. This process does not alter the overall configuration of the system. This can occur at any point, contributing a constant term to the energy. At the fourth order, bosons hop along a link to a neighboring site and then return along the same link, contributing a constant energy. Such processes are present for all links. This process does not change the positions of the bosons and, from the perspective of the dual lattice, the local constraint of two dimers per site remains satisfied. At the sixth order, one process involves a single boson hopping around the triangular face of each tetrahedron. Another process, as shown in Fig. 2(b), features atoms alternating between the ground and excited states along a hexagonal cell. This alternating configuration can be mapped, via a ring-exchange process, onto a complementary and flippable configuration in the dimer model. This process does not violate the local constraint (two dimers per site). The former contributes a constant energy. Ignoring all these constant terms, the effective Hamiltonian is a six-site ring exchange term
$H_{\textrm{eff}}^{(6)}=-t\sum\left(S_{1}^{+}S_{2}^{-}S_{3}^{+}S_{4}^{-}S_{5}^{+}S_{6}^{-}+h.c.\right),$ (4)
@@
where the sum is over all the hexagonal cells in the pyrochlore lattices. $t$ can be obtained by summing over virtual processes and $S_{i}^{\pm}$ represent that boson (spin) creation or annihilation at a single site $i$. Although this effective Hamiltonian is derived in the specific case of n=2, it is worth noting that for n=1 and n=3, the effective Hamiltonian we obtain is all described by the same universal form, namely Eq. (4).
The Schrieffer-Wolff perturbation theory in the context of quantum spin ice has also described in the paper [58]. This perturbation theory holds as long as the low-energy manifold hosts fixed number of Rydberg atoms within each tetrahedra (ice rule). For finite systems, as long as the tetrahedra structures are complete at the boundaries, the Rydberg blockade mechanism guarantees the low-energy configurations to obey such ice rule. As a result, the Schrieffer-Wolff formalism still applies and our results are applicable to finite-size spin systems given that tetrahedra structures are complete along the boundaries.
@@
For mapping to a QDM on the dual lattice, we connect the centers of all nearest-neighbor tetrahedra as shown in Fig. 1 (b) and 2 (b), with each atom exactly at the center of each link. Atoms in the Rydberg state (spin up) can be mapped to dimers in the dual (diamond) lattice, while atoms in the ground state (spin down) can be mapped to empty links. The high-order ring exchange dynamics becomes the effective kinetic term. Thus the QDM on diamond lattice takes the role of low-energy effective description of this pyrochlore Rydberg array [58]. The ground state of quantum dimer models can be exactly solved by adding a potential energy term $V$ and adjusting it to a particular value $V=t$, referred to as the RK point [40]. At the microscopic level, this $V$ term corresponds to high-order density interactions between Rydberg atoms and is not realistic in Rydberg experiments. However, in the context of pyrochlore quantum spin ice, it is reported in [67] that the third-nearest-neighbor Ising interaction $J_{3}$ plays a similar role to $V$, that can tune the system to the field-theoretical RK point with vanishing photon velocity. In Rydberg platform, this $J_{3}$ term also presents as the third-nearest-neighbor van der Waals interactions between Rydberg atoms.
@@
where the sum is performed over all hexagonal plaquettes. The first term is the ring exchange term and the second term is potential energy term.
@@
When $t/V=1$, i.e., at the famous RK point [40], the ground state is a resonating valence bond (RVB) state described by the coherent equal-weight superposition of all the dimer covering configurations, i.e., $|G\rangle=\sum_{C}|C\rangle$. Moreover, the RK point exhibits topological degeneracy, where the equal-weight superpositions within different topological sectors are ground states with identical energies [68].
@@
Figure 1: The mapping between spin, dimer and flux. (a) Rydberg array can be mapped to the spin model. The atom at the Rydberg excited state is mapped to the spin-up (up arrows), while the atom at the atomic ground state is mapped to the spin-down (down arrows).(b) Spin can be mapped onto dimer (orange line). A dimer, corresponds to a spin pointing upward, $S_{z}=1/2$, while the absence of a dimer (blue line) corresponds to a spin pointing downward, $S_{z}=-1/2$. (c) They can be interpreted as fluxes, we identify a spin pointing up, with a flux vector $B$ from sublattice a to b, and vice versa.
The ice rule corresponds to the contact of two upward spins, two dimers and the flux field vanishes for each tetrahedron.
@@
In the phase diagram of QDMs on 3D bipartite lattices, the RK point is located as a part of the U(1) Coulomb phase. In the spin language, the local constraint of dimers corresponds to the ice rule that the total spin in each tetrahedron of the pyrochlore lattice is same. In the dimer language, it means that the number of dimers attached to each site of the dual diamond lattice is fixed. Further, this local constraint is an essential ingredient that allows a mapping to the U(1) gauge theory: Spin-up (dimer) and spin-down (empty link) can be mapped to the $\pm\frac{1}{2}$ magnetic flux pointing from a to b diamond sublattices, as displayed in Fig. 1 [65, 69]. This leads to the condition $\left(\mathrm{div}B\right)_{\mathsf{r}}=0$ for the case of two-dimers-per-site configurations, and $\left(\mathrm{div}B\right)_{\mathsf{r}}=\eta_{\mathsf{r}}$ for the one-dimer-per-site configurations. Here $\mathsf{r}$ denotes the site of the dual diamond lattice, and $\eta_{\mathsf{r}}=\pm 1$ for a and b sublattices, respectively. Such divergence condition is analogous to Gauss’s law of electromagnetism, and the background charges can be absorbed by redefining the strength of the magnetic field.
@@
The pyrochlore lattice is a composite, consisting of four sets of face-centered cubic lattices. As shown in Fig. 2 (c), we label the points on the tetrahedron belonging to different sublattices as 1, 2, 3, and 4. In Fig. 2 (d), we display one of the sublattices and define its direct lattice vectors as $\overrightarrow{a_{1}},\overrightarrow{a_{2}},\overrightarrow{a_{3}}$, with the corresponding reciprocal lattice vectors as $\overrightarrow{b_{1}},\overrightarrow{b_{2}},\overrightarrow{b_{3}}$ (see Fig. 2 (e)).